OpenStudy (anonymous):

its can be shown that while deltax = dx, deltay is not neccesarily equal to dy. show this is true by comparing deltay and dy using the following. y=1-2x^2, x=2 and deltax=0.01 differentials i get the right answer for deltay=0.0401 but i get dy=-0.08 so i am lost

OpenStudy (anonymous):

\[y=1-2x^2\]First:\[y(2)=1-2*(2)^2=-7\\y(2+\Delta x)=y(2+0.01)=y(2.01)=1-2*(2.01)^2=-7.0802\]\[\Delta Y=y(2.01)-y(2)=-7.0802-(-7)=0.0802\]Now we have to do the derivative of y(x)\[y'(x)=-4x\]\[y'(2)=-8\\y'(2.01)=-8.04\]\[dy=y'(2.01)-y'(2)=-8.04-(-8)=0.04\]

OpenStudy (anonymous):

umm, shouldnt the answers be negative?

OpenStudy (anonymous):

ohh, sorry, my mistake.\[dy=y'(2.01)-y'(2)=-8.04-(-8)=-0.04\]

OpenStudy (anonymous):

same thing with

OpenStudy (anonymous):

\[DeltaY\]

OpenStudy (anonymous):

\[\Delta Y=-0.0802\]

OpenStudy (anonymous):

thaank you that cleared up that thank you